1. Do this problem by using only PW method: What is the IRR of a 20-year annual cash-flow investment of $1020 per year if the first year amount of $1020 is invested at the end of the first year, and the investment pays a single payment of $66,000 at the end of the 20th year?
Answer #1

Let I% be the IRR, then

66000 * (P/F, I%, 20) - 1020 * (P/A, I%,20) = 0

using trail and error method

When I = 10%, value of 66000 * (P/F, I%, 20) - 1020 * (P/A, I%,20) = 66000 * 0.148644 - 1020 * 8.513564 = 1126.6444

When I = 11%, value of 66000 * (P/F, I%, 20) - 1020 * (P/A, I%,20) = 66000 * 0.124034 - 1020 * 7.963328 = 63.6432

When I = 12%, value of 66000 * (P/F, I%, 20) - 1020 * (P/A, I%,20) = 66000 * 0.103667 - 1020 * 7.469444 = -776.8260

using interpolation

I = 11% + (63.6432 - 0) / (63.6432 + 776.8260) *(12% - 11%)

I = 11% + 0.0756%

I = 11.07% (Approx)